Stock R has a beta of 1.6, Stock S has a beta of 0.9, the required return on an average stock is 10%, and the risk-free rate of return is 7%. By how much does the required return on the riskier stock exceed the required return on the less risky stock? Round your answer to two decimal places.

%
Answer #1

required return= risk-free rate +Beta*(market rate- risk-free rate )

required return for R=7+1.6*(10-7)=11.8%

required return for S=7+0.9*(10-7)=9.7%

Hence excess required return=11.8%-9.7%

=2.1%